any oxidation-reduction reactions can be balanced by i spection. Try to balance the following reactions by inspection In each reaction, identify the substance reduced and the sub stance oxidized. a. Al(s) + HCl(aq)→AICI3(aq) 12(g) b. CH4(g) + S(s) → CS2(l) + H2S(g) c. C3H8(g) + O2(g) → CO2(g) + H2O(1) d. Cu(s) + Ag+(aq) → Ag(s) + Cu2+ (aq)
Answer #1

Balanced equation:
2 Al(s) + 6 HCl(aq) ====> 2 AlCl3(aq) + 3 H2(g)

Reaction type: single replacement

Aluminium is oxidized and H is reduced

Balanced equation:
4 S(s) + CH4(g) = CS2(l) + 2 H2S(g)

Reaction type: double replacement

Sulfur is reduced and Carbon is oxidized

Balanced equation:
C3H8(g) + 5 O2(g) = 3 CO2(g) + 4 H2O(l)

Reaction type: double replacement

Carbon is oxidized and oxygen is reduced

Balanced equation:
Cu(s) + Ag{+}(aq) = Ag(s) + Cu{+}(aq)

Reaction type: single replacement

Copper is oxidized and Silver is reduced